# An Excited Dark Matter Solution to the MeV Galactic Center Excesses

### v1

@@ 3 Results
We compare the 511 keV flux from Eq. (6) with INTEGRAL/SPI data, which provide fluxes binned in Galactic longitude $l$ and latitude $b$. Following Siegert_2023, we construct a longitude profile by binning in $3^{\circ}$ intervals in $l$ and integrating over $10.75^{\circ}<b<10.25^{\circ}$, and a latitude profile by binning in $3^{\circ}$ intervals in $b$ and integrating over $5.25^{\circ}<l<3.75^{\circ}$. For a given XDM model, we define
@@
Figure 3: 511 keV longitude (left) and latitude (right) profiles for XDM with $m_{\chi}=1.5$ TeV and $\Delta m=4$ MeV, compared to INTEGRAL/SPI data (red points), for three propagation setups: “Default” (solid), “Fast” (dashed) and “Slow” (dash-dotted), following DelaTorreLuque:2023cef.
@@
The bottom panels of Fig. 2 show the longitude and latitude profiles for fixed $\Delta m=4$ MeV and varying $m_{\chi}$. This choice of splitting produces monoenergetic positrons with $E_{e^{\pm}}=\Delta m/2=2$ MeV, which is the best-fit injection energy for the bulge component of the IfA spectrum in Knodlseder:2025pnx. For $m_{\chi}=1.5$ TeV and $\Delta m=4$ MeV we obtain an excellent fit to the SPI morphology with a reduced chi-squared of $\chi_{\mathrm{red}}^{2}\simeq 1.5$ for an NFW profile, even without invoking the IfA or CMZ constraints.
@@
At low $m_{\chi}$ the predicted bulge emission peaks sharply around the GC and can still be made compatible with SPI data once a disk component from astrophysical positron sources is included, such as massive stars or SNe, following Cappiello:2023qwl; laTorreLuquePedro:2024est. Such a disk component is guaranteed at some level, although its modelling carries significant uncertainties. Conversely, a cuspier DM profile naturally yields a cuspier 511 keV morphology and allows larger $m_{\chi}$ to remain compatible with SPI. This is illustrated in Fig. 6, where we compare the NFW case to a contracted NFW (cNFW) profile with inner slope $\gamma=1.2$. In this case the density profile, the velocity dispersion and escape velocity (Fig. 1), and thus the radial dependence of $\langle\sigma v\rangle$ from Eq. (5), are all modified.
@@
Fig. 3 shows the impact of CR propagation uncertainties on the 511 keV morphology. We consider the “Fast”, “Default” and “Slow” models of DelaTorreLuque:2023cef, which bracket optimistic and pessimistic choices for the diffusion and energy-loss parameters. The resulting differences in the longitude and latitude profiles are modest, reflecting the low kinetic energy of the injected positrons: their transport is dominated by energy losses rather than spatial diffusion. The main conclusions regarding the XDM fit to SPI are therefore robust against reasonable variations of the propagation setup.
@@
Using the best-fit cross sections $\sigma_{\mathrm{mr}}$ obtained from the 511 keV morphology for the models in Fig. 2, we compute the corresponding IfA spectra in the COMPTEL energy range. The top panels of Fig. 4 compare the all-sky XDM-induced emission to the bulge excess extracted by Knodlseder:2025pnx. Internal bremsstrahlung (IB) associated with the outgoing $e^{\pm}$ is included following Eq. (2) of Beacom:2004pe, with $m_{\chi}$ replaced by $\Delta m/2$, the positron energy. The separate IB and IfA components for our fiducial model are shown in the bottom panel.
@@
while the best-fit value reported by Knodlseder:2025pnx for a conventional 2.1 MeV annihilating DM scenario (including IB) is $\Phi_{511}=1.4\times 10^{-4}\ \mathrm{ph\,s^{-1}\,cm^{-2}}$. Given the systematic uncertainties inherent to the 511 keV measurement and the modelling of the IfA signal, this level of agreement indicates that the XDM framework passes a non-trivial consistency check: the values of $\sigma_{\mathrm{mr}}$ required to fit the SPI morphology also predict an MeV continuum compatible with the COMPTEL excess.
@@
Fig. 5 shows the ionization rate $\zeta(R)$ in the CMZ obtained from Eq. (10) using the propagated XDM-induced $e^{\pm}$ fluxes from DRAGON2. For a fixed splitting $\Delta m=4$ MeV and an NFW profile, the predicted ionization rates reach a few $\times 10^{-16}$ s-1 inside the CMZ, above standard CR expectations and remaining relatively flat with radius, broadly consistent with the inferred profile (Oka_2005; Oka_2020; Ravikularaman_2025) but still below the observed $\zeta_{\mathrm{H}_{2}}\gtrsim 10^{-15}\,\mathrm{s^{-1}}$.
@@
Obolentseva_2024 argued that ionization rates inferred for many molecular clouds across the Galaxy were biased high due to overestimates of their line-of-sight sizes and densities. This bias cannot by itself account for the extreme ionization levels in the CMZ, whose geometry and gas properties are better constrained, but it may lower the inferred CMZ ionization somewhat. Allowing for such a downward revision, the XDM-induced contribution could lie closer to the true ionization level and plausibly account for a non-negligible fraction of the excess. A more dedicated treatment of CR and DM transport in the complex environment of the CMZ is needed to sharpen this conclusion, but our results already show that the same XDM parameter space that fits the SPI and COMPTEL data naturally produces a significant, radially flat ionization component in the CMZ.
4 Discussion and Conclusions
@@
Using DRAGON2, we performed the first full CR propagation calculation of positrons in the XDM context. For $\Delta m\simeq 3$66 MeV and TeV-scale $m_{\chi}$, the resulting 511 keV longitude and latitude profiles are in strong agreement with SPI data, while the corresponding IfA spectrum matches the COMPTEL MeV bump within current statistical and systematic uncertainties. The same parameter space yields CMZ ionization rates $\zeta\sim 10^{-16}$$10^{-15}\,\mathrm{s^{-1}}$, larger than expected from standard CRs and with an approximately flat radial profile, although still below the nominal level inferred from molecular tracers. These late-time excitation rates can arise naturally from Sommerfeld enhancement in a light-mediator model and remain compatible with unitarity, thermal freeze-out and BBN constraints along the lines discussed in Finkbeiner:2007kk.
A light boson of mass $\sim 10$–20 MeV coupling to $e^{\pm}$ can play the role of the mediator in such XDM scenarios. Intriguingly, this mass range overlaps with the mild ($\sim 2\sigma$) excess near $16.9$ MeV reported by the PADME experiment PADME:2025dla, suggesting that forthcoming fixed-target searches may begin to probe exactly the mediator parameter space relevant for such Galactic MeV signatures. In parallel, future “MeV gap” missions such as COSI Karwin_2023; Siegert_2020 will provide sharper line and continuum measurements and improved morphology, offering a direct test of the correlated line and continuum predictions of XDM.
Beyond the Milky Way, massive galaxy clusters provide another promising arena. Their high velocity dispersions make inelastic upscattering efficient throughout the halo, potentially leading to non-thermal X-ray or $\gamma$-ray emission and small contributions to intracluster-medium heating. Current limits on diffuse cluster emission already require any such heating to be subdominant to active galactic nucleus feedback, but the possibility that DM–DM scattering produces modest intracluster heating or mild SZ distortions remains open and may be probed by upcoming high-sensitivity X-ray and SZ surveys SimonsObservatory:2018koc; CMB-S4:2016ple.
@@
5 Acknowledgements
SB is supported by the STFC under grant ST/X000753/1. DC acknowledges support from a Science and Technology Facilities Council (STFC) Doctoral Training Grant. PDL is supported by the Juan de la Cierva JDC2022-048916-I grant, funded by MCIU/AEI/10.13039/501100011033 European Union ”NextGenerationEU”/PRTR. The work of PDL is also supported by the grants PID2021-125331NB-I00 and CEX2020-001007-S, both funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. PDL also acknowledges the MultiDark Network, ref. RED2022-134411-T. This project used computing resources from the National Academic Infrastructure for Supercomputing in Sweden (NAISS) under project NAISS 2024/5-666. This project used computing resources from King’s College London CREATE system KCLCREATE2025. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising from this submission.

### v2

@@ III Results
We compare the 511 keV flux from Eq. (6) with INTEGRAL/SPI data, which provide fluxes binned in Galactic longitude $l$ and latitude $b$. Following Siegert (2023), we construct a longitude profile by binning in $3^{\circ}$ intervals in $l$ and integrating over $10.75^{\circ}<b<10.25^{\circ}$, and a latitude profile by binning in $3^{\circ}$ intervals in $b$ and integrating over $5.25^{\circ}<l<3.75^{\circ}$. For a given XDM model, we define
@@
Figure 3: 511 keV longitude (left) and latitude (right) profiles for XDM with $m_{\chi}=1.5$ TeV and $\Delta m=4$ MeV, compared to INTEGRAL/SPI data (red points), for three propagation setups: “Default” (solid), “Fast” (dashed) and “Slow” (dash-dotted), following De la Torre Luque et al. (2024).
@@
The bottom panels of Fig. 2 show the longitude and latitude profiles for fixed $\Delta m=4$ MeV and varying $m_{\chi}$. This choice of splitting produces monoenergetic positrons with $E_{e^{\pm}}=\Delta m/2=2$ MeV, which is the best-fit injection energy for the bulge component of the IfA spectrum in Knödlseder et al. (2025). For $m_{\chi}=1.5$ TeV and $\Delta m=4$ MeV we obtain an excellent fit to the SPI morphology with a reduced chi-squared of $\chi_{\mathrm{red}}^{2}\simeq 1.5$ for an NFW profile, even without invoking the IfA or CMZ constraints.
@@
At low $m_{\chi}$ the predicted bulge emission peaks sharply around the GC and can still be made compatible with SPI data once a disk component from astrophysical positron sources is included, such as massive stars or SNe, following Cappiello et al. (2023); De la Torre Luque; Pedro et al. (2025). Such a disk component is guaranteed at some level, although its modelling carries significant uncertainties. Conversely, a cuspier DM profile naturally yields a cuspier 511 keV morphology and allows larger $m_{\chi}$ to remain compatible with SPI. This is illustrated in Fig. 6, where we compare the NFW case to a contracted NFW (cNFW) profile with inner slope $\gamma=1.2$. In this case the density profile, the velocity dispersion and escape velocity (Fig. 1), and thus the radial dependence of $\langle\sigma v\rangle$ from Eq. (5), are all modified.
@@
Fig. 3 shows the impact of CR propagation uncertainties on the 511 keV morphology. We consider the “Fast”, “Default” and “Slow” models of De la Torre Luque et al. (2024), which bracket optimistic and pessimistic choices for the diffusion and energy-loss parameters. The resulting differences in the longitude and latitude profiles are modest, reflecting the low kinetic energy of the injected positrons: their transport is dominated by energy losses rather than spatial diffusion. The main conclusions regarding the XDM fit to SPI are therefore robust against reasonable variations of the propagation setup.
@@
Using the best-fit cross sections $\sigma_{\mathrm{mr}}$ obtained from the 511 keV morphology for the models in Fig. 2, we compute the corresponding IfA spectra in the COMPTEL energy range. The top panels of Fig. 4 compare the all-sky XDM-induced emission to the bulge excess extracted by Knödlseder et al. (2025). Internal bremsstrahlung (IB) associated with the outgoing $e^{\pm}$ is included following Eq. (2) of Beacom et al. (2005), with $m_{\chi}$ replaced by $\Delta m/2$, the positron energy. The separate IB and IfA components for our fiducial model are shown in the bottom panel.
@@
while the best-fit value reported by Knödlseder et al. (2025) for a conventional 2.1 MeV annihilating DM scenario (including IB) is $\Phi_{511}=1.4\times 10^{-4}\ \mathrm{ph\,s^{-1}\,cm^{-2}}$. Given the systematic uncertainties inherent to the 511 keV measurement and the modelling of the IfA signal, this level of agreement indicates that the XDM framework passes a non-trivial consistency check: the values of $\sigma_{\mathrm{mr}}$ required to fit the SPI morphology also predict an MeV continuum compatible with the COMPTEL excess.
@@
Fig. 5 shows the ionization rate $\zeta(R)$ in the CMZ obtained from Eq. (10) using the propagated XDM-induced $e^{\pm}$ fluxes from DRAGON2. For a fixed splitting $\Delta m=4$ MeV and an NFW profile, the predicted ionization rates reach a few $\times 10^{-16}$ s-1 inside the CMZ, above standard CR expectations and remaining relatively flat with radius, broadly consistent with the inferred profile (Oka et al., 2005; Oka and Geballe, 2020; Ravikularaman et al., 2025) but still below the observed $\zeta_{\mathrm{H}_{2}}\gtrsim 10^{-15}\,\mathrm{s^{-1}}$.
@@
Obolentseva et al. (2024) argued that ionization rates inferred for many molecular clouds across the Galaxy were biased high due to overestimates of their line-of-sight sizes and densities. This bias cannot by itself account for the extreme ionization levels in the CMZ, whose geometry and gas properties are better constrained, but it may lower the inferred CMZ ionization somewhat. Allowing for such a downward revision, the XDM-induced contribution could lie closer to the true ionization level and plausibly account for a non-negligible fraction of the excess. A more dedicated treatment of CR and DM transport in the complex environment of the CMZ is needed to sharpen this conclusion, but our results already show that the same XDM parameter space that fits the SPI and COMPTEL data naturally produces a significant, radially flat ionization component in the CMZ.
IV Discussion and Conclusions
@@
Using DRAGON2, we performed the first full CR propagation calculation of positrons in the XDM context. For $\Delta m\simeq 3$66 MeV and TeV-scale $m_{\chi}$, the resulting 511 keV longitude and latitude profiles are in strong agreement with SPI data, while the corresponding IfA spectrum matches the COMPTEL MeV bump within current statistical and systematic uncertainties. The same parameter space yields CMZ ionization rates $\zeta\sim 10^{-16}$$10^{-15}\,\mathrm{s^{-1}}$, larger than expected from standard CRs and with an approximately flat radial profile, although still below the nominal level inferred from molecular tracers. These late-time excitation rates can arise naturally from Sommerfeld enhancement in a light-mediator model and remain compatible with unitarity, thermal freeze-out and BBN constraints along the lines discussed in Finkbeiner and Weiner (2007).
A light boson of mass $\sim 10$–20 MeV coupling to $e^{\pm}$ can play the role of the mediator in such XDM scenarios. Intriguingly, this mass range overlaps with the mild ($\sim 2\sigma$) excess near $16.9$ MeV reported by the PADME experiment Bossi and others (2025), suggesting that forthcoming fixed-target searches may begin to probe exactly the mediator parameter space relevant for such Galactic MeV signatures. In parallel, future “MeV gap” missions such as COSI Karwin et al. (2023); Siegert et al. (2020) will provide sharper line and continuum measurements and improved morphology, offering a direct test of the correlated line and continuum predictions of XDM.
Beyond the Milky Way, massive galaxy clusters provide another promising arena. Their high velocity dispersions make inelastic upscattering efficient throughout the halo, potentially leading to non-thermal X-ray or $\gamma$-ray emission and small contributions to intracluster-medium heating. Current limits on diffuse cluster emission already require any such heating to be subdominant to active galactic nucleus feedback, but the possibility that DM–DM scattering produces modest intracluster heating or mild SZ distortions remains open and may be probed by upcoming high-sensitivity X-ray and SZ surveys Ade and others (2019); Abazajian and others (2016).
@@
V Acknowledgements
SB is supported by the STFC under grant ST/X000753/1. DC acknowledges support from a Science and Technology Facilities Council (STFC) Doctoral Training Grant. PDL is supported by the Juan de la Cierva JDC2022-048916-I grant, funded by MCIU/AEI/10.13039/501100011033 European Union ”NextGenerationEU”/PRTR. The work of PDL is also supported by the grants PID2021-125331NB-I00 and CEX2020-001007-S, both funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. PDL also acknowledges the MultiDark Network, ref. RED2022-134411-T. This project used computing resources from the National Academic Infrastructure for Supercomputing in Sweden (NAISS) under project NAISS 2024/5-666. This project used computing resources from King’s College London CREATE system King’s College London . For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising from this submission.